# DLinear-based Prediction of Remaining Useful Life of Lithium-Ion Batteries: Feature Engineering through Explainable Artificial Intelligence

### v1

@@ Abstract
Accurate prediction of the Remaining Useful Life (RUL) of lithium-ion batteries is essential for ensuring safety, reducing maintenance costs, and optimizing usage. However, predicting RUL is challenging due to the nonlinear characteristics of the degradation caused by complex chemical reactions. Machine learning allows precise predictions by learning the latent functions of degradation relationships based on cycling behavior. This study introduces an accurate RUL prediction approach based on feature engineering and DLinear, applied to the dataset from NASA’s Prognostics Center of Excellence. Among the 20 features generated from current, voltage, temperature, and time provided in this dataset, key features contributing to degradation are selected using Pearson correlation coefficient and Shapley values. Shapley value-based feature selection effectively reflects cell-to-cell variability, showing similar importance rankings across all cells. The DLinear-based RUL prediction using key features efficiently captures the time-series trend, demonstrating significantly better performance compared to Long Short-Term Memory and Transformer models.
@@
Keywords— Lithium-Ion Batteries, Remaining Useful Life, Feature Engineering, DLinear, Cell-to-Cell Variability
@@ 1 Introduction
The Electric Vehicles (EVs) industry has grown rapidly due to environmental issues related to internal combustion engine vehicles. Lithium-Ion Batteries (LIBs) are the most widely used battery materials based on their advantages such as high energy density, low self-discharge characteristics, and long lifespan [1]. However, it is inevitable that battery performance will degrade due to capacity fade and impedance increase that occurs during cycling. Monitoring capacity reduction and accurately predicting Remaining Useful Life (RUL) through battery management systems have become crucial, as the performance degradation of lithium-ion batteries can end up with critical risks such as thermal runaway reactions [2, 3]. Predicting the necessary replacement time in advance is challenging, due to intra-cell factors including morphology, architecture, and composition, as well as inter-cell factors such as winding, connection, and assembly processes that contribute to Cell-to-Cell Variability (CtCV) [4, 5], along with differences in operating conditions for each device. The Prognostics and Health Management (PHM) system implements condition-based management by predicting the necessary replacement timing before failures occur, allowing for repairment only when it’s necessary [6, 7]. It considers battery capacity degradation using various sensor data, such as voltage, current, and temperature. RUL prediction is one of the critical functions of the PHM system, which forecasts the remaining cycles where the battery can be used safely. Accurate RUL prediction contributes to optimizing battery replacement timing and prevents accidents caused by excessive charging or discharging.
@@
RUL prediction methods using early cycling behavior can generally be classified into two approaches: model-based approach and data-driven approach [8, 9]. However, model-based approaches often encounter challenges. In practice, the irreversible capacity fade is caused by an interaction of mechanical (e.g., particle cracking) and chemical (e,g., electrolyte decomposition) factors, making it quite difficult to mathematically identify these degradation mechanisms. Additionally, there can be a knee point where performance degradation, which progresses slowly in the early stages, accelerates after a certain cycle. This nonlinearity of performance degradation also contributes to the difficulty of model-based approaches. Machine Learning (ML)–based approach is a promising alternative to overcome this challenge [10, 11, 12].
@@
ML–based RUL prediction research typically employs time series algorithms, with changes observed for each charging/discharging cycle. Traditionally, Recurrent Neural Network (RNN) [13, 14, 15] and Long Short-Term Memory (LSTM) [16] have been used for such tasks: Liu et al. [17] propose an adaptive RNN that continuously updates the parameters of the network to improve the description performance of dynamic cycling behavior of the LIBs. Park et al. [18] conducted RUL forecasting using LSTM, but they struggle with learning long-term dependency [19]. To overcome this limitation, the attention algorithm was introduced, enabling predictions that consider a broader context rather than just neighboring time steps [20, 21]. Designed for natural language processing task and centered around the attention algorithm, Transformer architectures outperforms RNN and LSTM by taking a bird’s-eye view of time-series data to learn features of data, then trains in order of importance rather than sequential dependency [22]. Chen et al. [23] conducted a prediction of RUL by Transformer, successfully overcoming the long-term dependency hurdle of RNN and LSTM. Despite being a game changer in time series forecasting, the Transformer has encountered difficulties when applied to time series prediction. Specifically, its high complexity in space (i.e., memory) time and its performance are not significantly superior to conventional LSTM models.
An alternative, DLinear, was proposed to address these challenges [24]. DLinear uses two linear layers considering trend and seasonality, in contrast to the complex structure and extensive computation of Transformer. Unlike the Transformer, DLinear takes both trend and seasonality into account in time series data predictions. Furthermore, it provides explainablity with its weights of linear layers, revealing parameter importance. For battery life, RUL typically follows a decreasing trend with occasional increases. While the Transformer struggles to learn this trend, DLinear captures both the decreasing and periodic increasing trends, outperforming LSTM and Transformer-based predictions in aspects of space complexity and time complexity.
@@
The performance of ML–based RUL prediction such as DLinear is determined by the quantity and quality of data. Since battery cycling experiments require a significant budget in time and cost, there are practical difficulties in generating large-scale datasets. Therefore, it is important to identify features that significantly contribute to battery lifespan and achieve high prediction performance even with a small number of training cycles. However, because high-dimensional features can overfit the cycling behavior used for training, selecting key features is essential, and various feature engineering studies have been conducted to predict battery lifespan: Fei et al. [25] extracted 42 features from a large-scale lithium iron phosphate (LFP) battery dataset [11] and identified a high-performance prediction model structure by applying various feature selection approaches such as wrapper and filter approaches and various prediction models. Hu et al. [26] proposed a fusion approach that combines the advantages of wrapper and filter approaches to select features that contribute to state-of-health prediction. Greenbank et al. [27] propose an automatic feature selection method using the Pearson Correlation Coefficient (PCC), a filter approach, to predict the knee point where battery life declines rapidly and input the End-of-Life (EoL) prediction. However, the most widely used filter approach has concerns that correlation ranking can only consider linear dependencies between features and output. Shapley Additive exPlanation (SHAP) [28], which has recently been used in sensitivity analysis, provides insight at the feature selection stage because it presents qualitative results as well as quantitative sensitivity [29, 30]. Additionally, in Ref. [31], SHAP value-assisted feature selection was found to have superior performance compared to other algorithms for time series applications such as battery lifespan prediction. Moreover, the SHAP-based approach, which is a model-agnostic approach, is useful when applied to various machine learning models. In this study, we identify features that contribute to RUL prediction through the filter approach and SHAP-based approach and apply them to a ML model.
@@
Here, we identify the features that contribute to the battery lifespan and predict the RUL of the battery using the DLinear model. Features contributing to battery capacity degradation were extracted based on battery chemistry. For the entire feature set, each key feature subset identified through the filter approach and SHAP-based approach is used as input for LSTM, Transformer, and DLinear, and the accuracy and EoL cycle are compared. The rest of this paper is structured as follows. In Section 2, the three machine learning algorithms used for the prediction of RUL are described. In Section 3, the dataset used in the cycling experiment and the scheme for extracting features contributing to the battery lifespan based on battery chemistry are explained. Additionally, two algorithms used for feature selection are introduced. Section 4 discusses feature engineering and RUL prediction results, and Section 5 presents a summary of the results and future work.
2 Time-series prediction model
Figure 1: Schematic diagram of RUL prediction through feature engineering
@@ 2.1 Long short-term memory (LSTM)
Sequential Neural Network (SNN) process time-series data by assuming that the state at any given time step depends on prior time steps. RNN is the pioneering SNN structure designed for processing time-series data [13, 14, 15]. The RNN model takes input to produce output, and also sends its output back to itself, to recurrent into the next time step. Due to this structure, each neuron in RNN has two sets of weights: one for input and another for multiplying with the output from the previous time step. In this way, RNN takes previous information into account and determines the extent of this reference with weights. However, it has pointed that as time progresses, the continuous multiplication of weights in RNN results in the dilution of information from distant past time steps at the current time step, a problem known as gradient vanishing [32].
@@ 2.2 Transformer
For a long time, LSTM has been firmly established as the state-of-the-art approach in time series data. However, long-term state of LSTM also did not completely solve the gradient vanishing problem [19]. The problem was alleviated but similar, as the consideration of information from distant past time steps being diluted as weights were multiplied with the long-term state of LSTM with the progression of time steps. Over time, an Attention algorithm was devised to focus more on contextually significant information after receiving information, not only from the immediate previous time step but from all time steps [20, 21]. This allowed for learning based on meaningful information from the entire dataset rather than focusing solely on nearby information, thereby addressing the issues of RNN, LSTM, and similar SNN algorithms. Furthermore, implemented with the Attention algorithm, the Transformer was devised. Recently, Transformer-based approaches and variant models of Transformer were utilized in predicting the RUL of LIBs [23].
@@ 2.3 DLinear
The Transformer has successfully overcome the hurdle of gradient vanishing and long-range dependencies posed by LSTM. However, as the Transformer was designed for natural language processing tasks (e.g., translation, summarization), not for time-series data, it had the weakness of poor performance when applied to time-series data. In particular, early Transformer models had difficulty in identifying trends in data that increased or decreased continuously throughout the observation period. Furthermore, due to its complex structure for analyzing sentences and texts written in languages, the Transformer required excessive computation compared to the conventional SNNs, thereby demanding a significant amount of time and memory. To address these two issues, DLinear was proposed [24].
Algorithm 1 Decomposition-Linear (DLinear) Model for RUL Prediction
1:Input: Feature matrix $X:[F_{1},F_{2},...,F_{20}]$, Target $y=\text{RUL}$
@@
5:if $individual$ is True then
@@
The detailed procedure for RUL prediction using DLinear is introduced in Algorithm 1. DLinear separates time-series data into Trend ($X_{t}$) and Remainder ($X_{s}$) using a moving average, and learns each through a separate linear layer. Equation of $X_{t}$ and $X_{s}$ can be defined as follows:
@@ 3 Feature engineering
In this section, feature engineering to extract parameters contributing to RUL from cycling data of LIBs and identify key parameters is explained. Our RUL prediction framework is validated using the NASA dataset [33]. In Section 3.1, the process of extracting features expected to contribute to lifespan through information on voltage, temperature, and time for four LIBs cells (i.e., B0005, B0006, B0007, and B0018) is described. In Section 3.2, the process by which a significantly contributing subset of the entire feature set is selected through the filter approach and SHAP-based approach is explained.
@@ 3.1 Feature extraction
In feature extraction, factors that contribute to RUL prediction are extracted based on battery cycling data collected from the Battery Management System (BMS). According to battery chemistry, the main factors that accelerate degradation are voltage [34] and temperature [35] during cycling, and degradation causes differences in charge and discharge profiles. Table 1 describes 20 features categorized by current and voltage, temperature, and time. Fig. 2(a-c) shows that the current and voltage profile at the charging stage, temperature distribution, and discharge voltage curve are dependent on the RUL. These dependent characteristics are processed through statistical information such as skewness and included in the feature set. In particular, the characteristic of the discharge curve, where differences occur after significant cycles, is explained by the slope at the early stage of the discharge (Fig. 2(d)).
@@
Current and Voltage F1 Variance of measured voltage in discharging
F2 Variance of measured current in discharging
@@
Figure 2: Feature description of charging and discharging for B0005 battery cell: (a) CC charging time (F18), (b) Temperature in discharging, (c) Voltage in discharging, (d): Slope of discharge voltage curve (F6, F7, and F8)
@@ 3.2.1 Filter approach
The filter approach uses feature ranking to select key features. In general, the correlation between the extracted features and the objective function (i.e., RUL) scores the ranking of each feature, selects features that are higher than a certain threshold, and screens other features [36]. Because the filter approach is applied without being integrated into the prediction model, it can be easily applied to high-dimensional feature sets with relatively low computational budgets. This quantifies the linear correlation between the feature denoted by $x$ and the objective function denoted by $y$.
@@
Where $\bar{x}$ and $\bar{y}$ are sample mean of $x$ and $y$, respectively. As for the key feature selection, all correlation coefficients between key features and cell capacity are evaluated based on Eq. 3. The three parameter subsets with the highest absolute correlation coefficients are regarded as the most sensitive features, while those with the lowest absolute values are considered the least sensitive features. This concept also applies to the SHAP-based approach described in Section 3.2.2

### v2

# State-of-Health Prediction for EV Lithium-Ion Batteries via DLinear and Robust Explainable Feature Selection

@@ Abstract
Accurate prediction of the state-of-health (SOH) of lithium-ion batteries is essential for ensuring the safety, reliability, and efficient operation of electric vehicles (EVs). Battery packs in EVs experience nonuniform degradation due to cell-to-cell variability (CtCV), posing a major challenge for real-time battery management. In this work, we propose an explainable, data-driven SOH prediction framework tailored for EV battery management systems (BMS). The approach combines robust feature engineering with a DLinear. Using NASA’s battery aging dataset, we extract twenty meaningful features from voltage, current, temperature, and time profiles, and select key features using Pearson correlation and Shapley additive explanations (SHAP). The SHAP-based selection yields consistent feature importance across multiple cells, effectively capturing CtCV. The DLinear algorithm outperforms long short-term memory (LSTM) and Transformer architectures in prediction accuracy, while requiring fewer training cycles and lower computational cost. This work offers a scalable and interpretable framework for SOH forecasting, enabling practical implementation in EV BMS and promoting safer, more efficient electric mobility.
@@
Keywords— Lithium-ion batteries, State-of-health, Feature engineering, DLinear, Cell-to-cell variability
@@ 1 Introduction
The electric vehicles (EVs) industry has grown rapidly due to environmental issues related to internal combustion engine vehicles. Lithium-ion batteries (LIBs) are the most widely used battery materials based on their advantages such as high energy density, low self-discharge characteristics, and long lifespan [1, 2, 3]. However, it is inevitable that battery performance degrades as impedance increases due to chemical decomposition and lithium plating during battery usage. Monitoring capacity reduction and accurately predicting state-of-health (SOH) through battery management system (BMS) have become crucial, as the performance degradation of LIBs can pose critical risks, such as thermal runaway reactions [4, 5, 6]. Predicting the necessary replacement time in advance is challenging, due to intra-cell factors including morphology, architecture, and composition, as well as inter-cell factors such as winding, connection, and assembly processes that contribute to cell-to-cell variability (CtCV) [7, 8], along with differences in operating conditions for each device. The prognostics and health management (PHM) system implements condition-based management by predicting the necessary replacement chance before failures occur, allowing for repairment only when it’s necessary [9, 10]. It considers battery capacity degradation using various sensor data, such as voltage, current, and temperature. SOH prediction is one of the critical functions of the PHM system, which forecasts the remaining cycles where the battery can be used safely. Accurate SOH prediction contributes to optimizing battery replacement timing and prevents accidents caused by excessive charging or discharging.
@@
SOH prediction methods using early cycling behavior can generally be classified into two approaches: model-based approach and data-driven approach [11, 12]. However, model-based approaches often encounter challenges. In practice, the irreversible capacity fade is caused by an interaction of mechanical (e.g., particle cracking) and chemical (e.g., electrolyte decomposition) factors, making it quite difficult to mathematically identify these degradation mechanisms. The existence of a knee point at which aging accelerates in a certain cycle, and often a slight capacity increase early in the cycle, makes model-based approaches difficult [13]. Furthermore, while battery models such as the Doyle-Fuller-Newman (DFN) [14, 15] model have been successfully applied to predict SOH, the strong nonlinearity of these models necessitates further analysis, including identifiability analysis [16, 17]. The data-driven approach is a promising alternative to overcome this challenge [18, 13, 19].
@@
The data-driven SOH prediction research typically employs time-series machine learning (ML) algorithms, with changes observed for each charging/discharging cycle. Normally, recurrent neural network (RNN) [20, 21, 22] and long short-term memory (LSTM) [23] have been used for such tasks: Liu et al. [24] propose an adaptive RNN that continuously updates the parameters of the network to improve the description performance of dynamic cycling behavior of the LIBs. Park et al. [25] conducted SOH forecasting using LSTM, but they struggle with learning long-term dependency [26]. To overcome this limitation, the attention algorithm was introduced, enabling predictions that consider a broader context rather than just neighboring time steps [27, 28]. Designed for natural language processing tasks and centered around the attention algorithm, Transformer architectures outperform RNN and LSTM by taking a holistic view of time-series data to learn features of data, then trains in order of importance rather than sequential dependency [29]. Chen et al. [30] conducted a prediction of SOH by Transformer, successfully overcoming the long-term dependency hurdle of RNN and LSTM. Despite being a game-changer in time series forecasting, the Transformer has encountered difficulties when applied to time series prediction [31]. Specifically, its high complexity in space (i.e., memory) time and its performance are not significantly superior to conventional LSTM models. To address these challenges, DLinear was proposed [31]. DLinear uses two linear layers considering trend and seasonality, in contrast to the complex structure and extensive computation of the Transformer. Unlike the Transformer, DLinear takes both trend and seasonality into account in time series data predictions. Furthermore, it provides explainablity with its weights of linear layers, revealing parameter importance. The SOH of the battery typically exhibits a decreasing pattern during its lifespan, with capacity increases occasionally observed at the early cycles, punctuated by periodic increases. Although the Transformer finds it challenging to grasp this trend, DLinear effectively captures both the downward and occasional upward movements, surpassing LSTM and Transformer-based predictions in terms of both space and time complexity.
@@
The performance of data-driven SOH prediction, such as DLinear, is determined by the quantity and quality of data. Since battery cycling experiments require a significant budget in time and cost, there are practical difficulties in generating large-scale datasets. Therefore, it is essential to identify the key features that significantly contribute to battery health and to attain high predictive performance, even with a limited number of training cycles. However, because high-dimensional features can overfit the cycling behavior used for training, selecting key features is essential, and various feature engineering studies have been conducted to predict battery lifespan: Fei et al. [32] extracted 42 features from a large-scale lithium iron phosphate (LFP) battery dataset [13] and identified a high-performance prediction model structure by applying various feature selection approaches such as wrapper and filter approaches and various prediction algorithm. Hu et al. [33] proposed a fusion approach that combines the advantages of wrapper and filter approaches to select features that contribute to SOH prediction. Greenbank et al. [34] propose an automatic feature selection method using the Pearson correlation coefficient (PCC), a filter approach, to predict the knee point where battery life declines rapidly and input the end-of-life (EoL) prediction. However, the most widely used filter approach has concerns that correlation ranking can only consider linear dependencies between features and output. Shapley additive explanation (SHAP) [35], which has recently been used in sensitivity analysis, provides insight at the feature selection stage because it presents qualitative results as well as quantitative sensitivity [36, 37]. Additionally, in Ref. [38], SHAP value-assisted feature selection was found to have superior performance compared to other algorithms for time series applications such as battery lifespan prediction. Moreover, the SHAP-based approach, which is a model-agnostic approach, is useful when applied to various data-driven models. Based on the above considerations, this work focuses on feature engineering based on battery chemistry and data-driven prediction of SOH. Through sensitivity analysis, key features identified are used as inputs to a time-series forecasting model to accurately predict the decreasing trend of SOH. Specifically, our main contributions are presented below:
  1. 1)
    Using the data collected from the cycling experiment, a feature set required for SOH prediction is generated. PCC and SHAP-based global sensitivity analysis are performed to identify key features among all features.
  2. 2)
    The key feature subset is used as input to LSTM, Transformer, and DLinear, and the accuracy is compared with previous works.
  3. 3)
    The variability of cells composing modules or packs is intensively analyzed. Key features that are equally applied to the cells within specific modules or packs provide opportunities for efficient SOH prediction in practical BMS.
@@
This article is organized as follows: In Section 2, the three machine learning algorithms used for the prediction of SOH are described. In Section 3, the dataset used in the cycling experiment and the scheme for extracting features contributing to the battery lifespan based on battery chemistry are explained. Additionally, two algorithms used for feature selection are introduced. Section 4 discusses feature engineering and SOH prediction results, and Section 5 presents a summary of the results and future work.
2 Time-series prediction algorithm
Figure 1: Schematic diagram of SOH prediction through feature engineering
@@ 2.1 Long short-term memory (LSTM)
Sequential neural network (SNN) process time-series data by assuming that the state at any given time step depends on prior time steps. RNN is the pioneering SNN structure designed for processing time-series data [20, 21, 22]. The RNN algorithm takes input to produce output, and also sends its output back to itself, to recurrent into the next time step. Due to this structure, each neuron in RNN has two sets of weights: one for input and another for multiplying with the output from the previous time step. In this way, RNN takes previous information into account and determines the extent of this reference with weights. However, it has pointed that as time progresses, the continuous multiplication of weights in RNN results in the dilution of information from distant past time steps at the current time step, a problem known as gradient vanishing [39].
@@ 2.2 Transformer
For a long time, LSTM has been firmly established as the state-of-the-art approach in time series data. However, long-term state of LSTM also did not completely solve the gradient vanishing problem [26]. The problem was alleviated but similar, as the consideration of information from distant past time steps being diluted as weights were multiplied with the long-term state of LSTM with the progression of time steps. Over time, an Attention algorithm was devised to focus more on contextually significant information after receiving information, not only from the immediate previous time step but from all time steps [27, 28]. This allowed for learning based on meaningful information from the entire dataset rather than focusing solely on nearby information, thereby addressing the issues of RNN, LSTM, and similar SNN algorithms. Furthermore, implemented with the Attention algorithm, the Transformer was devised. Recently, Transformer-based approaches and variant models of Transformer were utilized in predicting the SOH of LIBs [30].
@@ 2.3 DLinear
The Transformer has successfully overcome the hurdle of gradient vanishing and long-range dependencies posed by LSTM. However, as the Transformer was designed for natural language processing tasks (e.g., translation, summarization), not for time-series data, it had the weakness of poor performance when applied to time-series data. In particular, early Transformer models had difficulty in identifying trends in data that increased or decreased continuously throughout the observation period. Furthermore, due to its complex structure for analyzing sentences and texts written in languages, the Transformer required excessive computation compared to the conventional SNNs, thereby demanding a significant amount of time and memory. The DLinear model was introduced to mitigate these two problems [31]. Its inherently low computational burden, stemming from the linear structure of the model, offers a distinct advantage for BMS, as these systems necessitate real-time forecasting and are often constrained by the practical limitations of deploying models with excessive computational demands.
Algorithm 1 Decomposition-Linear (DLinear) Algorithm for SOH Prediction
1:Input: Feature matrix $X:[F_{1},F_{2},...,F_{20}]$, Target $y=\text{SoH}$
@@
5:if individualindividual is True then
@@
The detailed procedure for SOH prediction using DLinear is introduced in Algorithm 1. DLinear separates time-series data into Trend ($X_{t}$) and Remainder ($X_{s}$) using a moving average, and learns each through a separate linear layer. Equation of $X_{t}$ and $X_{s}$ can be defined as follows:
@@ 3 Feature engineering
In this section, feature engineering to extract parameters contributing to SOH from cycling data of LIBs and identify key parameters is explained. Our SOH prediction framework is validated using the LIBs aging dataset provided by NASA’s Prognostics Center of Excellence [40]. The nominal capacity of the four selected cells (i.e., B0005, B0006, B0007, and B0018) is 2 Ah. All the cells are cycled under the same charge and discharge procedure. In the charge phase, the cells are charged with a constant current (CC) of 1.5A until the voltage reaches 4.2V. After the voltage reaches 4.2V, they are charged with a constant voltege (CV) until the current reaches 0.02A. In the discharge phase, the cells are discharged in CC mode with 2.0 A until B0005, B0006, B0007, and B0018 reach 2.7V, 2.5V, 2.2V, and 2.5V, respectively. In Section 3.1, the process of extracting features expected to contribute to lifespan through information on voltage, temperature, and time for four lithium-ion batteries (LIBs) cells is described. In Section 3.2, the process by which a significantly contributing subset of the entire feature set is selected through the filter approach and SHAP-based approach is explained.
@@ 3.1 Feature extraction
In feature extraction, factors that contribute to SOH prediction are extracted based on battery cycling data collected from the BMS. According to battery chemistry, the main factors that accelerate degradation are voltage [41] and temperature [42] during cycling, and degradation causes differences in charge and discharge profiles. Consequently, even if a feature that cannot be explained by battery chemistry is identified as a key feature [13], it is still necessary to find key features that can explain a battery’s health for SOH estimation. Table 1 describes 20 features generated from current, voltage, temperature, and time data collected in each cycle. Fig. 2(a-c) shows that the current and voltage profile at the charging stage, temperature distribution, and discharge voltage curve are dependent on the SOH. These dependent characteristics are processed through statistical information, such as skewness, and included in the feature set. In particular, the characteristic of the discharge curve, where differences occur after significant cycles, is explained by the slope at the early stage of the discharge (Fig. 2(d)).
@@
Current and Voltage F1 Variance of measured current in discharging
F2 Variance of measured voltage in discharging
@@
Figure 2: Feature description of charging and discharging for B0005 battery cell: (a) CC charging time (F18), (b) Temperature in discharging, (c) Voltage in discharging, (d) Slope of discharge voltage curve (F6, F7, and F8)
@@ 3.2.1 Filter approach
The filter approach uses feature ranking to select key features. In general, the correlation between the extracted features and the objective function (i.e., SOH) scores the ranking of each feature, selects features that are higher than a certain threshold, and screens other features [43]. Because the filter approach is applied without being integrated into the prediction model, it can be easily applied to high-dimensional feature sets with relatively low computational budgets. This quantifies the linear correlation between the feature denoted by $x$ and the objective function denoted by $y$.
@@
Where $\bar{x}$ and $\bar{y}$ are sample mean of $x$ and $y$, respectively. As for the key feature selection, all correlation coefficients between key features and cell capacity are evaluated based on Eq. (3). The three parameter subsets with the highest absolute correlation coefficients are regarded as the most sensitive features, while those with the lowest absolute values are considered the least sensitive features. This concept also applies to the SHAP-based approach described in Section 3.2.2